Design of Algorithms - Course Review

K. Subramani LCSEE, West Virginia University, Morgantown, WV {ksmani@csee.wvu.edu}

March 31, 2011

1. Enumerated below are various notions of teaching effectiveness. Kindly pick five that closely match your notion of effectiveness and rank them accordingly. Also feel free to add some of your own criterial, but remember that only five can be ranked.

(a) Arriving to class in timely fashion.

(b) High enthusiasm in instruction.

(c) Uniform, fair and consistent grading.

(d) A good syllabus which is consistent with course expectations in peer schools.

(e) High expectations from instructor of students.

(f) Ability to inculcate critical thinking.

(g) Tests and exams which emphasize deep thinking as opposed to learning by rote.

(h) Intellectually engaging the students.

(i) If my grades meet my expectations, the instructor is effective.

(j) Should keep the students satisfied.

(k) Course and lecture organization.

(l) Good feedback/communication to outline work assignments.

(m) Providing lots of help to students.

(n) Providing minimum help to students with the goal of encouraging independent thinking.

(o) Good articulation of material.

(p) Avoids repetition of content.

(q) Avoids criticizing students in order to not offend them, although this could jeopardize the student’s learning.